# Adaptive Gradient Calibration for Single-Positive Multi-Label Learning in Remote Sensing Image Scene Classification

### v1

@@ Abstract
Multi-label classification (MLC) offers a more comprehensive semantic understanding of Remote Sensing (RS) imagery compared to traditional single-label classification (SLC). However, obtaining complete annotations for MLC is particularly challenging due to the complexity and high cost of the labeling process. As a practical alternative, single-positive multi-label learning (SPML) has emerged, where each image is annotated with only one relevant label, and the model is expected to recover the full set of labels. While scalable, SPML introduces significant supervision ambiguity, demanding specialized solutions for model training. Although various SPML methods have been proposed in the computer vision domain, research in the RS context remains limited. To bridge this gap, we propose Adaptive Gradient Calibration (AdaGC), a novel and generalizable SPML framework tailored to RS imagery. AdaGC adopts a gradient calibration (GC) mechanism combined with Mixup and a dual exponential moving average (EMA) module for robust pseudo-label generation. To maximize AdaGC’s effectiveness, we introduce a simple yet theoretically grounded indicator to adaptively trigger GC after an initial warm-up stage based on training dynamics, thereby guaranteeing the effectiveness of GC in mitigating overfitting to label noise. Extensive experiments on two benchmark RS datasets under two distinct label noise types demonstrate that AdaGC achieves state-of-the-art (SOTA) performance while maintaining strong robustness across diverse settings.
@@ I Introduction
With the rapid development of satellite missions, vast amounts of remote sensing (RS) imagery have become increasingly accessible, serving as a critical data source for numerous applications [1, 2]. As a fundamental task, image classification plays a vital role in connecting imagery with downstream applications in the RS domain [3, 4, 5]. However, most existing RS studies adopt the single-label classification (SLC) paradigm, assigning a single semantic label to each image [6, 7], which oversimplifies the inherent complexity of real-world RS scenes [8, 9]. In practice, RS images usually contain diverse land cover types, objects, or functional zones within a single scene. As illustrated in Fig. 1, very-high resolution (VHR) imagery often captures land use and land cover information as well as fine-grained objects such as buildings, vehicles, and vegetation within a limited area. Lower-resolution imagery often spans heterogeneous regions, such as agricultural mosaics or urban–rural transitions. In both cases, multi-label classification (MLC) offers a more informative, comprehensive, and realistic semantic representation of RS imagery compared to traditional SLC methods.
However, MLC faces a major challenge in acquiring complete and accurate annotations due to the difficulty and high cost in the annotation process [10]. This challenge is particularly critical in the deep learning (DL) era, where large-scale labelled datasets are essential [11]. A practical compromise is to annotate each image with only one relevant label, even when multiple classes are present, which leads to the single-positive multi-label learning (SPML) setting [12]. In this setting, the model is weakly supervised with one positive label per image, yet it is expected to recover the complete set of relevant labels. Supervision in SPML is inherently weak, characterized by missing labels. In practice, even expert annotators often miss relevant labels in multi-label settings, making the false negative issue pervasive in MLC. SPML can thus be viewed as an extreme case of this problem and also offers a promising solution to completing label sets in single-label annotation scenarios.
@@
Figure 1: Single- and multi-label annotation examples from (a) AID-multilabel [13] and (b) refined BigEarthNet [14] datasets, where the corresponding CLC mask is presented for reference. Compared to the single-class labels, the multi-label annotations can more comprehensively describe the scene.
@@
Although SPML is promising in scalability and annotation efficiency, it also introduces significant ambiguity that can bias the learning process. A common baseline approach is the AN strategy [12], which treats all unobserved labels as negatives, unavoidably introducing bias into learning. In the computer vision (CV) domain, several strategies have been proposed to address its limitations, such as noise-robust regularization terms, online label-correction/pseudo-labeling methods, and interpretability-driven frameworks [12, 15, 16, 17, 18, 19, 20]. Despite the advancements in CV, the SPML problem remains underexplored in RS. SPML is particularly valuable for RS image classification, where annotation typically requires greater domain expertise and incurs higher costs compared to natural image datasets. These challenges make the application of SPML in RS more necessary and also more difficult.
@@
When directly applying existing SPML methods to RS data, several challenges arise. One major issue is the reliability of pseudo-labeling methods. VHR RS images often include small, sparsely distributed targets, while medium-to-low-resolution imagery tends to capture broader regional semantics. Both cases can easily result in low inter-class separability and high intra-class variability, hindering the effectiveness of simple pseudo-labeling strategies such as confidence thresholding. Furthermore, the distinct characteristics of MLC versus multi-class classification (MCC) suggest that commonly used learning-from-noisy-labels strategies, which are often designed for MCC, should be carefully tailored to SPML. Moreover, although early learning theory indicates that models gradually reach a peak generalization performance in the early stage of training before overfitting to label noise [21, 22], most existing SPML methods rely on predefined warm-up schedules and overlook this dynamic behaviour. This, to some extent, limits the model’s capability in pseudo-labeling.
@@
In this work, we propose a novel and generalizable framework for SPML in RS, termed Adaptive Gradient Calibration (AdaGC), to address the aforementioned challenges. The framework consists of three key components. First, we introduce a dual exponential moving average (EMA) module for robust pseudo-label estimation, which leverages the consensus between the temporally smoothed predictions of the student model via EMA and the outputs of the teacher model, whose weights are updated via EMA of the student’s parameters. Second, the generated pseudo-labels are used to guide a Gradient Calibration (GC) mechanism inspired by Early Learning Regularization (ELR) [21], which is originally designed for MCC, to alleviate overfitting to false negatives. To further improve generalizability, we incorporate the effective data augmentation technique Mixup [15] within the GC process. Finally, to ensure the effectiveness of this calibration process, we conduct an in-depth analysis of learning dynamics in SPML and propose a simple yet theoretically grounded indicator for adaptive activation of GC after an initial warm-up stage. This mechanism is crucial for avoiding the negative influence of low-quality pseudo-labels on model optimization. We evaluate the AdaGC on two benchmark RS multi-label datasets with two single-positive label simulation ways. Extensive experiments demonstrate the effectiveness and robustness of AdaGC.
@@
  • We propose a two-stage SPML framework for RS imagery based on a student–teacher architecture, incorporating an early learning detection strategy that guides the activation of GC to combat overfitting to label noise. Furthermore, we provide theoretical analyses that support the effectiveness of our approach in SPML scenarios.
@@
  • We perform extensive empirical evaluations on two benchmark RS datasets, covering high- and low-resolution imagery, and two single positive label noise simulation strategies (Random and Dominant). We also provide a comprehensive benchmarking of existing SPML methods, offering valuable insights for future research in this domain.
@@ II Related Works
This section reviews advancements in MLC, first focusing on its application in RS and then discussing in general the challenges posed by noisy labels. Finally, we delve into SPML, detailing various methods developed in CV to address its unique difficulties.
@@ II-A Remote Sensing Image Multi-label Classification
Early DL research in MLC for RS imagery primarily focused on high-resolution data and employed Convolutional Neural Networks for feature extraction, followed by classifiers such as Radial Basis Function Neural Networks [23] or Structured Support Vector Machines [24]. These approaches often relied on conventional transfer learning, using CNNs pre-trained on generic CV datasets (e.g., ImageNet [25]) as fixed feature extractors. However, the significant differences between natural CV scenes and complex RS imagery can lead to suboptimal classification performance, limiting the effectiveness of such direct transfer.
To mitigate the limitations of pre-trained models, alternative strategies have been proposed. For instance, Stivaktakis et al. [26] introduced a data augmentation scheme to facilitate end-to-end training of shallow CNNs, along with replacing the softmax layer with a sigmoid function to accommodate multi-label outputs. While this direct adaptation remains a common practice in multi-label scenarios, it may result in the imprecise identification of multiple classes. More sophisticated approaches integrate sequential neural network architectures with CNNs to enhance multi-label scene classification. Hua et al. [8] proposed a class-wise attention-based Recurrent Neural Network (RNN) to explicitly model the co-occurrence relationships among multiple classes, thus generating class predictions sequentially. Similarly, an attention-aware label relational reasoning network was introduced in [13] to localize discriminative regions and characterize inter-label relationships based on feature maps. Alshehri et al. [27] presented an encoder-decoder neural network architecture, where the encoder includes a squeeze-and-excitation layer to model channel-wise dependencies, and the RNN-based decoder serves as an adaptive spatial attention module. These attention-driven strategies have demonstrated considerable success in identifying informative areas through attention maps derived from convolutional features.
@@ II-B Multi-label Classification with Noisy Labels
A major hurdle in MLC in RS is the difficulty in obtaining exhaustive multi-label annotations, which often results in noisy labels. This noise can be categorized as either missing labels (subtractive noise), where a class present in the considered image is not annotated, or as wrong labels (additive noise), where an absent class is incorrectly annotated [10]. Such multi-label noise, which can be introduced through zero-cost annotation (e.g., using thematic products), can severely impair model training. Consequently, developing methods robust to multi-label noise in RS is an active research area, usually categorized as weakly supervised multi-label learning (WSML) [28].
Methods for handling noisy labels in multi-label settings often draw inspiration from related fields, such as positive-unlabeled learning (PUL) [29, 30, 31] and noisy multi-class learning [32]. Common strategies involve developing robust loss functions that are less sensitive to label noise or class imbalance [33]. Another approach is pseudo-labeling, where likely labels for unobserved or unknown annotations are inferred to provide additional training signals [34]. The “memorization effect,” where models tend to learn clean labels early but later start to memorize noisy ones, has been widely observed in noisy MCC. ELR [21] was developed to specifically counteract this, by adjusting the influence of gradients from potentially noisy samples, particularly in the early stages of training. Burgert et al. [10] later extended it to MLC in a straightforward manner and empirically demonstrated its effectiveness. Additional strategies include sample selection, which aims to identify and potentially exclude samples with noisy labels [35], and label correction/refurbishment, which attempts to correct the noisy labels (e.g., using collaborative learning frameworks [36, 37]).
Other methods, such as those in the large loss (LL) [16] family, including LL-R, LL-Ct, and LL-Cp, are developed to recognize and handle mislabeled samples with large losses, based on the observation that noisy labels often lead to high loss values. Specifically, LL-R rejects samples with large losses, LL-Ct applies temporary corrections, and LL-Cp employs permanent corrections. BoostLU [19] complements LL by enhancing predicted positive regions via a function applied to class activation maps, further improving the robustness.
@@ III Methodology
Figure 2: Flowchart of the proposed Adaptive Gradient Calibration (AdaGC) method for single positive multi-label learning in remote sensing image classification.
As illustrated in Fig. 2, the proposed method adopts a student–teacher framework, where the teacher model’s weights are updated via EMA. The training pipeline consists of two stages. In the first stage, the student model is trained directly on noisy labels, during which the teacher model’s validation accuracy (with respect to noisy labels) is monitored as an early learning indicator to determine when to activate the second stage— Gradient Calibration (GC). Next, we first define the problem before presenting the technical details of the proposed method.
@@ III-B Gradient Calibration
The concept of gradient calibration for handling noisy labels was first proposed in [21] for MCC, where ELR introduces an auxiliary regularization term to constrain the standard cross-entropy loss, thus preventing the memorization of noisy labels. The regularization term for MCC is defined as follows:
$\mathcal{R}_{\text{MCC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\log\left(1-\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle\right),$ (3)
where $n$ is the number of training samples, $\mathbf{p}$ denotes the softmax output vector of the multi-class model, and $\mathbf{t}$ represents the soft pseudo-labels. The gradient of (3) is defined as follows:
@@
$\mathbf{g}_{i,c}=\frac{p_{i,c}}{1-\langle\mathbf{p}_{i},\mathbf{t}_{i}\rangle}\sum_{k=1}^{C}\left(t_{i,k}-t_{i,c}\right)\cdot p_{i,k}.$ (5)
@@
$\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\log\left(1-p_{i,c}\cdot t_{i,c}\right).$ (7)
We can compute the gradient of (7) as follows:
$\nabla_{\theta}\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\nabla_{\theta}f(\mathbf{x}_{i};\theta)\cdot g_{i,c},$ (8)
@@
$g_{i,c}=\frac{p_{i,c}-1}{\frac{1}{p_{i,c}\cdot t_{i,c}}-1}\leq 0.$ (9)
We argue that (7) is particularly suitable for the SPML setting, as it specifically penalizes false negatives. Accordingly, we apply GC only to negative labels, excluding its use for positive ones in practice. The final objective is a combination of the binary cross-entropy loss and the GC regularization:
@@
However, applying GC from the beginning may lead the model to predict all classes as positive. Liu et al. [22] show that adaptively triggering correction before overfitting to label noise is crucial for the effectiveness of sample-correction-related techniques. Inspired by this, in Section III-C we introduce a simple early learning detection mechanism to appropriately activate GC. Moreover, as GC performance depends on the quality of the pseudo-labels $\mathbf{t}$, in Section III-D we design a dual-EMA module to ensure the robust generation of pseudo-labels.

### v2

@@ Abstract
Multi-label classification (MLC) offers a more comprehensive semantic understanding of Remote Sensing (RS) imagery compared to traditional single-label classification (SLC). However, obtaining complete annotations for MLC is particularly challenging due to the complexity and high cost of the labeling process. As a practical alternative, single-positive multi-label learning (SPML) has emerged, where each image is annotated with only one relevant label, and the model is expected to recover the full set of labels. While scalable, SPML introduces significant supervision ambiguity, demanding specialized solutions for model training. Although various SPML methods have been proposed in the computer vision domain, research in the RS context remains limited. To bridge this gap, we propose Adaptive Gradient Calibration (AdaGC), a novel and generalizable SPML framework tailored to RS imagery. AdaGC adopts a gradient calibration (GC) mechanism with a dual exponential moving average (EMA) module for robust pseudo-label generation. We introduce a theoretically grounded, training-dynamics-based indicator to adaptively trigger GC, which ensures GC’s effectiveness by preventing it from being affected by model underfitting or overfitting to label noise. Extensive experiments on two benchmark RS datasets under two distinct label noise types demonstrate that AdaGC achieves state-of-the-art (SOTA) performance while maintaining strong robustness across diverse settings. The codes and data will be released at https://github.com/rslab-unitrento/AdaGC.
@@ I Introduction
With the rapid development of satellite missions, vast amounts of remote sensing (RS) imagery have become increasingly accessible, serving as a critical data source for numerous applications [zhao_artificial_2024, zhu_foundations_2024]. Image classification [mehmood_remote_2022, dimitrovski_current_2023, he_recent_2018], a fundamental task to connect RS imagery with downstream applications, still largely adopts the single-label classification (SLC) paradigm. However, assigning a single label to each image oversimplifies the complexity of real-world RS scenes [Hua_2019_JPRS, Sumbul_2020_ACCESS]. In practice, RS images often contain diverse land cover types, objects, or functional zones. As illustrated in Fig. 1, both VHR and lower-resolution imagery commonly capture heterogeneous content, making multi-label classification (MLC) a more informative representation than SLC.
A key challenge for MLC is the difficulty and cost of acquiring complete, accurate annotations [Burgert_2022_TGRS], especially in the deep learning (DL) era where large-scale labeled datasets are crucial [zhu_deep_2017]. A practical workaround is to annotate only one relevant label per image, leading to the single-positive multi-label learning (SPML) setting [Cole_2021_CVPR]. In this setting, the model is weakly supervised with one positive label per image, yet it is expected to recover the complete set of relevant labels. The supervision in SPML is characterized by missing labels. In practice, even expert annotators often miss relevant labels, making the false negative issue pervasive in MLC. SPML can thus be viewed as an extreme case of this problem.
@@
Figure 1: Single- and multi-label annotation examples from (a) AID-multilabel [hua_relation_2020] and (b) refined BigEarthNet [clasen_reben_2025_arxivonly] datasets, where the corresponding CLC mask is presented for reference. Compared to the single-class labels, the multi-label annotations can more comprehensively describe the scene.
@@
Although SPML is promising in scalability and annotation efficiency, it introduces significant ambiguity. A common baseline approach is the AN strategy [Cole_2021_CVPR], which treats all unobserved labels as negatives, unavoidably introducing bias into learning. In the computer vision (CV) domain, several strategies have been proposed to address its limitations, such as noise-robust regularization terms, online label-correction/pseudo-labeling methods, and interpretability-driven frameworks [Cole_2021_CVPR, Zhang_2018_ICLR, Kim_2022_CVPR, Xie_2022_NIPS, Chen_2024_IJCAI, Kim_2023_CVPR, Liu_2023_ICML]. Despite the advancements in CV, the SPML problem remains underexplored in RS. SPML is particularly valuable for RS image classification, where annotation typically requires greater domain expertise and incurs higher costs. These challenges make the application of SPML in RS more necessary and also more difficult.
@@
When directly applying existing SPML methods to RS data, several challenges arise. One major issue is the reliability of pseudo-labels. very-high resolution (VHR) RS images often include small, sparsely distributed targets, while medium-to-low-resolution imagery tends to capture broader regional semantics. Both cases can easily result in low inter-class separability and high intra-class variability, hindering the effectiveness of simple pseudo-labeling strategies such as confidence thresholding. Furthermore, the distinct characteristics of MLC versus multi-class classification (MCC) suggest that commonly used learning-from-noisy-labels strategies, which are often designed for MCC, should be carefully tailored to SPML. Moreover, most existing pseudo-label-based SPML methods depend on fixed warm-up schedules, which often result in suboptimal performance due to model underfitting or overfitting to label noise.
@@
In this work, we propose a novel and generalizable framework for SPML in RS, termed Adaptive Gradient Calibration (AdaGC), to address the aforementioned challenges. The framework consists of three key components. First, we propose a theoretically grounded Gradient Calibration (GC) mechanism designed specifically for SPML. GC aims to alleviate the model’s tendency to overfit false negatives by correcting gradients using pseudo-label guidance. Unlike Early Learning Regularization (ELR) [liu_early-learning_2020-2], which is formulated for multi-class classification under the simplex constraint, GC is rigorously extended to the SPML setting without such a constraint. Second, to guarantee pseudo-label quality, we introduce a dual exponential moving average (EMA) module for robust pseudo-label estimation, which leverages the consensus between the temporally smoothed predictions of the student model and the teacher model’s outputs. The former applies EMA to predictions, and the latter to model weights. Then, to ensure the effectiveness of GC, we conduct an in-depth analysis of learning dynamics in SPML and propose a theoretically grounded indicator for adaptive activation of Gradient Calibration (GC) after an initial warm-up stage. This mechanism is central to our design, which is capable of preventing interference from underfitting or noise-driven overfitting to stabilize optimization. We also incorporate the Mixup data augmentation technique [Zhang_2018_ICLR] within GC to further improve model’s generalizability. Extensive experiments on two benchmark RS multi-label datasets demonstrate the effectiveness and robustness of AdaGC.
@@
  • We propose a two-stage RS SPML framework, which incorporates an early learning detection strategy to guide the activation of GC to combat underfitting and overfitting to label noise. The theoretical analyses are also provided to support the effectiveness of the framework design.
@@
  • We conduct extensive experiments on two benchmark RS datasets covering both high- and low-resolution imagery. For low-resolution data, we employ Random and Dominant SPML noise simulations to better reflect real-world annotation behavior. For high-resolution imagery, where identifying a dominant class is inherently difficult, we rely on Random simulation supplemented with additional Manual single-positive annotations. We provide a comprehensive benchmarking of existing SPML methods, offering valuable insights for future research in this domain.
@@ II Related Works
This section reviews advancements in MLC, with emphasis on RS applications, challenges from noisy labels, and recent SPML methods developed in CV.
@@ II-A Remote Sensing Image Multi-label Classification
Early DL research in MLC for RS imagery primarily focused on high-resolution data and employed Convolutional Neural Networks for feature extraction, followed by classifiers such as Radial Basis Function Neural Networks [Zeggada_2017_GRSL] or Structured Support Vector Machines [Koda_2018_TGRS]. These approaches often relied on conventional transfer learning, using CNNs pre-trained on generic CV datasets (e.g., ImageNet [Russakovsky_2015_IJCV]) as fixed feature extractors. However, the significant differences between natural CV scenes and complex RS imagery can lead to suboptimal classification performance, limiting the effectiveness of such direct transfer.
To mitigate the limitations of pre-trained models, alternative strategies have been proposed. For instance, Stivaktakis et al. [Stivaktakis_2019_GRSL] introduced a data augmentation scheme to facilitate end-to-end training of shallow CNNs, along with replacing the softmax layer with a sigmoid function to accommodate multi-label outputs. While this direct adaptation remains a common practice in multi-label scenarios, it may result in the imprecise identification of multiple classes. More sophisticated approaches integrate sequential neural network architectures with CNNs to enhance multi-label scene classification. Hua et al. [Hua_2019_JPRS] proposed a class-wise attention-based Recurrent Neural Network (RNN) to explicitly model the co-occurrence relationships among multiple classes, thus generating class predictions sequentially. Similarly, an attention-aware label relational reasoning network was introduced in [hua_relation_2020] to localize discriminative regions and characterize inter-label relationships based on feature maps. Alshehri et al. [Alshehri_2019_ACCESS] presented an encoder-decoder neural network architecture, where the encoder includes a squeeze-and-excitation layer to model channel-wise dependencies, and the RNN-based decoder serves as an adaptive spatial attention module. These attention-driven strategies have demonstrated considerable success in identifying informative areas through attention maps derived from convolutional features.
@@ II-B Multi-label Classification with Noisy Labels
A major hurdle in MLC in RS is the difficulty in obtaining exhaustive multi-label annotations, which often results in noisy labels. This noise can be categorized as either missing labels (subtractive noise), where a class present in the considered image is not annotated, or as wrong labels (additive noise), where an absent class is incorrectly annotated [Burgert_2022_TGRS]. Such multi-label noise, which can be introduced through zero-cost annotation (e.g., using thematic products), can severely impair model training. Consequently, developing methods robust to multi-label noise in RS is an active research area, usually categorized as weakly supervised multi-label learning (WSML) [Xie_2023_TPAMI].
Methods for handling noisy labels in multi-label settings often draw inspiration from related fields, such as positive-unlabeled learning (PUL) [duPlessis_2014_NIPS, Gang_2016_NIPS, bekker_2020_ML] and noisy multi-class learning [perantonirobust]. Common strategies involve developing robust loss functions that are less sensitive to label noise or class imbalance [li2020robust]. Another approach is pseudo-labeling, where likely labels for unobserved or unknown annotations are inferred to provide additional training signals [Zhang_2020_RS]. The “memorization effect,” where models tend to learn clean labels early but later start to memorize noisy ones, has been widely observed in noisy MCC. Early Learning Regularization (ELR) [Liu_2020_NIPS] was developed to specifically counteract this, by adjusting the influence of gradients from potentially noisy samples, particularly in the early stages of training. Burgert et al. [Burgert_2022_TGRS] later extended it to MLC in a straightforward manner and empirically demonstrated its effectiveness. Additional strategies include sample selection, which aims to identify and potentially exclude noisy labels [Paris_2021_TGRS], label correction/refurbishment, which attempts to correct noisy labels (e.g., using collaborative learning frameworks [Aksoy_2021_ICIP, Aksoy_2024_TNNLS]), and noise-robust regularization, such as using word embedding similarities to regularize model training [hua_learning_2020].
Other methods, such as those in the large loss (LL) [Kim_2022_CVPR] family, including LL-R, LL-Ct, and LL-Cp, are developed to recognize and handle mislabeled samples with large losses, based on the observation that noisy labels often lead to high loss values. Specifically, LL-R rejects samples with large losses, LL-Ct applies temporary corrections, and LL-Cp employs permanent corrections. BoostLU [Kim_2023_CVPR] complements LL by enhancing predicted positive regions via a function applied to class activation maps, further improving the robustness.
@@ III Methodology
Figure 2: Flowchart of the proposed Adaptive Gradient Calibration (AdaGC) method for single-positive multi-label learning in remote sensing image classification. Pseudo-labels $\mathbf{t}$ are generated by combining the teacher model’s predictions $\mathbf{p}^{T}$ and the student model’s predictions $\tilde{\mathbf{p}}^{S}$ according to (20). EMA is also applied to the student model’s predictions during the warm-up stage, which is omitted for simplification.
As illustrated in Fig. 2, the proposed method adopts a student–teacher architecture, where the teacher model’s weights are updated via EMA. The training pipeline consists of two stages. In the first stage, the student model is trained directly on noisy labels, during which the teacher model’s validation accuracy (with respect to noisy labels) is monitored as an early learning indicator to determine when to activate the second stage— Gradient Calibration (GC). Next, we first define the problem before presenting the technical details of AdaGC.
@@ III-B Gradient Calibration
The concept of gradient calibration for handling noisy labels was first proposed in [Liu_2020_NIPS] for MCC, where ELR introduces an auxiliary regularization term to constrain the standard cross-entropy loss, thus preventing the memorization of noisy labels. The regularization term for MCC is defined as follows:
$\mathcal{R}_{\text{MCC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\log\left(1-\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle\right).$ (3)
Here, both $\mathbf{p}_{i}$ and $\mathbf{t}_{i}$ are probability vectors whose elements are non-negative and sum to 1. This simplex constraint ensures that their inner product $\left\langle\mathbf{p}_{i},\mathbf{t}_{i}\right\rangle$ lies within the range $[0,1]$. The gradient of (3) is defined as follows:
@@
$g_{i,c}=\frac{p_{i,c}}{1-\langle\mathbf{p}_{i},\mathbf{t}_{i}\rangle}\sum_{k=1}^{C}\left(t_{i,k}-t_{i,c}\right)\cdot p_{i,k}.$ (5)
@@
$\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\mathbb{I}(y_{i,c}=0)\log\left(1-p_{i,c}\cdot t_{i,c}\right),$ (7)
where $\mathbb{I}(\texttt{True})=1$ otherwise 0. Considering the SPML setting, GC is applied only to negative labels, excluding its use for positive ones in practice. We can compute the gradient of (7) as follows:
$\nabla_{\theta}\mathcal{R}_{\text{MLC}}^{\text{GC}}\left(\theta\right)=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\mathbb{I}(y_{i,c}=0)\nabla_{\theta}f_{c}(\mathbf{x}_{i};\theta)\cdot g_{i,c},$ (8)
@@
$g_{i,c}=\frac{-t_{i,c}}{1-p_{i,c}\cdot t_{i,c}}\cdot{p_{i,c}(1-p_{i,c})}\leq 0,$ (9)
where $p_{i,c}(1-p_{i,c})$ is the sigmoid derivative. We argue that (7) is particularly suitable for the SPML setting, as it specifically penalizes false negatives by pushing the model to predict positive scores for (negative) labels with higher pseudo-label scores $t_{i,c}$. The final objective is a combination of the binary cross-entropy loss and the GC regularization:
@@
However, applying GC from the beginning may lead the model to predict all classes as positive. Liu et al. [liu_aio2_2024] show that adaptively triggering correction before overfitting to label noise is crucial for the effectiveness of sample-correction-related techniques. Inspired by this, in Section III-C we introduce a simple early learning detection mechanism to appropriately activate GC. Moreover, as GC performance depends on the quality of the pseudo-labels $\mathbf{t}$, in Section III-D we design a dual-EMA module to ensure the robust generation of pseudo-labels.